A mass of m is sliding on a inclined plane of mass M which is kept on smooth surface find displacement of mass M.
1
Expert's answer
2020-09-23T11:46:29-0400

let the angle of inclination be "\theta"


As clearly seen from the figure,

then the net force acting on mass m in downward direction is

f= "mgsin\theta-f_s"

Where "f_s" is frictional force on inclined surface

According to Newton's Law Third law,

This force also equal to the force on the inclined plane in upward direction as shown in figure


f="mgsin\theta-u_smgcos\theta"

Let the horizontal force on Mass M be F

then f=F cos"\theta"

"Fcos\theta=mg(sin\theta-u_scos\theta)"


"F=mg(tan\theta-u_s)" ....(1)


Let the acelaration of complete system be a

( m +M ) a=F

a="\frac{F}{(m+M)}"


Displacement in time t is given by second equation of motion as

s="ut+\frac{1}{2}at^2"

u=0(initial velocity is 0)

="\frac{1}{2}\times\frac{F}{m+M}\times t^2"

="\frac{mg(tan\theta-u_s)t^2}{2(m+M)}"